OpenStudy (anonymous):

(2x^2+4)(3x+3) over 7x-6 Find the derivative using quotient rule Help please!

hero (hero):

If I were you I would multiply that out first, but let me see something first, hold on....

OpenStudy (anonymous):

yeah multiply out in the numerator first will make your life much easier. you do not want to use the quotient rule combined with the product rule if you can avoid it

OpenStudy (anonymous):

i will be quiet now

OpenStudy (anonymous):

f'(x)=\[=((2x^2+4)(3x+3))'(7x-6)-(2x^2+4)(3x+3)(7x-6)' \over (7x-6)\]\[=((2x^2+4)'(3x+3)+(2x^2+4)(3x+3)')(7x-6)-(2x^2+4)(3x+3)(7x-6)' \over (7x-6)\]

OpenStudy (anonymous):

Sorry. Over (7x-6)^2

hero (hero):

Wow, you're really taking the long way on this one

OpenStudy (anonymous):

Thank you. Saved me hella typing.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

You smell. There. I said it. Now satellite doesn't have to.

OpenStudy (anonymous):

\[(2x^2+4)(3x+3)=6 x^3+6 x^2+12 x+12\]

hero (hero):

I hate this shiz

OpenStudy (anonymous):

sorry, just thought i would mention it

OpenStudy (anonymous):

i got that im not sure about my final anwer

OpenStudy (anonymous):

your first step should look like \[\frac{(7x-6)(18x^2+12x+12)-(6x^3+6x^2+12x+12)7}{(7x-6)^2}\] then a raft of annoying ugly algebra

OpenStudy (anonymous):

can yiu show me the rest..

hero (hero):

sprusty, don't let me catch you making a mistake, you will get the wrath from me.

OpenStudy (anonymous):

this one? do i really have to do this algebra in the numerator? it is multiply and combine like terms

OpenStudy (anonymous):

\[ \frac{6 (14 x^3-11 x^2-12 x-26)}{(7x-6)^2}\]there factored and all

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

can you help me on the other ones?